# FlipedRAG: Black-Box Opinion Manipulation Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 3.2 Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG system to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM, it calculates relevance score $R(q,d)$ for each document $d$ in $D$ and return the set of documents with the $k$ highest relevance score as the retrieval result (the LLM only takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)=\{d_{1},d_{2},...,d_{k}\}$. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$.
3.2.1 Objective of the Adversary
@@
The attacker aims to find the adversarial trigger $p_{\text{adv}}$ for each $d_{t}$, and adds it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $R(q,d_{t}\oplus p_{\text{adv}})$ assigned by the retrieval model RM, $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)=\{d\mid d_{t}\oplus p_{\text{adv}}\}$ , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\{d\mid d_{t}\oplus p_{\text{adv}}\}))=S_{t}$, $S_{t}\neq S_{o}$.
3.2.2 Capabilities of the Adversary
@@
In the black-box scenario, we assume the attacker is only authorized to query the RAG, obtain RAG-generated results and modify limited documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the attacker has no knowledge of the model architecture, model parameters, or any other information related to the models within the black-box RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited.
@@
Figure 3: The framework for obtaining ranking imitation data of retrieval model in black-box RAG system.
3.3 Black-box Retriever Imitation
@@
The primary issue in implementing manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to simulate the black-box retrieval model $RM$ in RAG. The basic idea is to train a surrogate model $M_{\text{i}}$ with the ranking results $RM_{\text{k}}(q)$ from the retrieval model $RM$, thus turning the black-box retrieval model into a white-box model. However, since the retriever and the LLM in RAG are serially connected, it is not possible to directly obtain the ranking results of the retriever. Therefore, we attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents irrelevant to the query can be random sampled as negative examples $d_{-}$ Therefore, we design specific instructions to make the black-box RAG replicate the retrieval results of the retriever RM. These retrieval results only need to reflect the relevance to the query. Then, based on the generated results of the LLM and the corpus, we sample positive and negative data to train the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG is illustrated in Figure 3. The specific instruction used is as follows:
@@
After that, we use a pairwise approach to sample data and train the surrogate model. The black-box RAG responds the instruction with context information so the responses generated by the black-box RAG reflect the retrieval results instead of being independent to the context. Relevant documents are sampled from the responses as positive examples $d_{+}$, and random irrelevant documents are sampled as negative examples $d_{-}$.
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Let the relevance score calculated by $M_{i}$ be $R_{i}$, we train the surrogate model $M_{i}$ by minimizing:
$L=-\frac{1}{|Q|}\sum_{q\in Q}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$       (1)
3.4 Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) [13] strategy for adversarial retrieval attacks, which is commonly used as baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus. Then, by querying the system of the black-box RAG model to get the generated response, we compare the opinions of the manipulated response with the opinions of the response generated by the benign RAG to evaluate the reliability of the black-box RAG.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its generated optimization function being [13]:
$\hskip 14.22636pt\max\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot\log P_{g}(T_{\text{pat}};w)\\ \hskip 71.13188pt+\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}\hskip 19.91692pt$ (1)
@@ 3.4 Opinion Manipulation Attack
Subsequently, we insert $p_{\text{adv}}$ into the target document to execute the poisoning attack on the RAG corpus. Our optimization objective of RAG opinion manipulation attack is as follows:
$\hskip 14.22636pt\max\dfrac{1}{|Q|}\sum_{q\in Q}\text{I}(\text{S}(LLM(q,RM_{k}(q)))=S_{t})\\ \hskip 14.22636pts.t.RM(q)=E(q)\cdot E(d),d\in D(d;d_{t}\oplus p_{\text{adv}})\hskip 56.9055pt\\ \hskip 14.22636ptRM_{k}(q)=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q)\}$ (2)
@@
4 Experiments
4.1 Dataset
In terms of dataset, we utilize the MS MARCO Passages Ranking dataset as the data source for guiding the black-box RAG to generate relevant passages [27] where we sample data pairs to train the surrogate model. Additionally, we use controversial topics data scraped from the PROCON.ORG website as the object of manipulation. The controversial topic dataset includes over 80 topics, covering fields such as society, health, government, and education. Each controversial topic is discussed from two stances (Pro and Con), with an average of 30 related passages, each holding a certain opinion with stance Pro or Con. We random select 30 public controversial topics to conduct opinion manipulation experiment on RAG.
4.2 Experiment Details
@@
(1) Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. Based on a retriever and a large language model (LLM), Conversational RAG mainly consists of a history-awaring retriever to identify users’ intention and a question-answering chain to produce output, as shown in Figure 4. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama3), Vicuna-13b-v1.5(Vicuna) and Mixtral-8x7B-Instruct-v0.1(Mixtral). They perform well across various tasks among all open-source LLMs. The prompt connecting the retriever and the LLM in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain:
Use the following pieces of retrieved context to answer the question. Keep the answer concise. Context: {context}. Question: {question}.
Figure 4: The framework of Conversational RAG.
@@
(2) Target retriever model and surrogate model. People usually use dense retrieval models as the retriever in RAG. Therefore, we select three representative dense retrieval models, Contriever, Co-Condenser and ANCE, as the target retrieval model [7, 8, 22]. By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset.
@@
(3) Manipulation target. For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial trigger $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3.
@@
(4)Imitation data construction. We collect imitation training data in a pairwise way, sampling both positive samples and negative samples. It is worth mentioning that we randomly sample negative examples from the top of the ranking produced by the original surrogate model. We discover that taking highly relevant documents ranked by the surrogate model as negative examples helps the surrogate model learn more features of the target black-box retrieval model by contrastive learning between positive and negative examples.
(5) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.

### v3

@@ 3.2. Threat Model
Given a topic $q$ (the question) from a set of controversial topics $Q$, a RAG corpus $D$ and a target opinion $S_{t}$ (Pro or Con, “Pro” stands for supporting and “Con” stands for opposing) as the opinion that the attacker expects the RAG system to generate, we denote the target document as $d_{t}$ which holds $S_{t}$ for $q$. Given a black-box retrieval model RM, it calculates relevance score $R(q,d)$ for each document $d$ in $D$ and returns the set of documents with the $k$ highest relevance score as the retrieval result (the LLM only takes top $k$ candidate documents into account). The retrieval result is denoted as $\text{RM}_{k}(q)=\{d_{1},d_{2},...,d_{k}\}$. The LLM of the black-box RAG utilizes $\text{RM}_{k}(q)$ to answer $q$ with response $\text{LLM}(q,RM_{k}(q))$, the opinion of $\text{LLM}(q,RM_{k}(q))$ is $S_{o}$, $S_{o}=\text{S}(LLM(q,RM_{k}(q)))$.
3.2.1. Objective of the Adversary
@@
The attacker aims to find the adversarial trigger $p_{\text{adv}}$ for each $N$ selected target document $d_{t}$, and add it to $d_{t}$, transforming the corpus $D$ to $D(d;d_{t}\oplus p_{\text{adv}})$. $p_{\text{adv}}$ can increase the relevance score $R(q,d_{t}\oplus p_{\text{adv}})$ assigned by the retrieval model RM, $d_{t}$ will be ranked at the top of the retrieval results $\text{RM}_{k}(q)=\{d\mid d_{t}\oplus p_{\text{adv}}\}$ , guiding the LLM to generate responses that align with the target opinion: $\text{S}(LLM(q,\{d\mid d_{t}\oplus p_{\text{adv}}\}))=S_{t}$, $S_{t}\neq S_{o}$.
3.2.2. Capabilities of the Adversary
@@
In the black-box scenario, we assume the attacker is only authorized to query the RAG, obtain RAG-generated results and modify at most $N$ documents in the corpus. There are no restrictions on the number of calls to RAG. Furthermore, the attacker has no knowledge of the model architecture, model parameters, or any other information related to the models within the black-box RAG. Modifying the prompt templates used by the LLM in RAG is also prohibited.
@@
Figure 3. The framework for obtaining ranking imitation data of retrieval model in black-box RAG system.
3.3. Black-box Retriever Imitation
@@
The primary issue in implementing manipulation is to make the retrieval model of the black-box RAG transparent so that the attacker is able to produce adversarial triggers based on the knowledge of black-box retriever. We aim to imitate the black-box retrieval model $RM$ in RAG. The basic idea is to train a surrogate model $M_{\text{i}}$ with the ranking results $RM_{\text{k}}(q)$ from the retrieval model $RM$, thus turning the black-box retrieval model into a white-box model. However, since the retriever and the LLM in RAG are serially connected, it is not possible to directly obtain the ranking results of the retriever. Therefore, we attempt to guide the LLM to replicate the retrieval results of the black-box RAG, then we obtain the documents deemed relevant to the user query by the black-box retrieval model $RM$, which can be used as positive examples $d_{+}$ for black-box imitation training. Subsequently, documents irrelevant to the query can be randomly sampled as negative examples $d_{-}$ Therefore, we design specific instructions to induce the black-box RAG to replicate the retrieval results of the retriever RM. These retrieval results only need to reflect the relevance to the query. Then, based on the generated results of the LLM and the corpus, we sample positive and negative data to train the surrogate model. The method for obtaining imitation data of the retrieval model in black-box RAG is illustrated in Figure 3. The specific instruction used is as follows:
@@
After that, we use a pairwise approach to sample data and train the surrogate model. The black-box RAG responds to the instruction with context information so the responses generated by the black-box RAG reflect the retrieval results. Relevant documents are sampled from the responses as positive examples $d_{+}$, and random irrelevant documents are sampled as negative examples $d_{-}$. For real-world RAG system attack, negative samples can be constructed using web information randomly sampled from search engines like Google to simulate real-world adversarial noise in RAG corpora.
These sample pairs $(d_{+},d_{-})$ are then incorporated into the training dataset. After sampling the imitation data, we employ a pairwise training method to train the surrogate model $M_{i}$. Let the relevance score calculated by $M_{i}$ be $R_{i}$, we train the surrogate model $M_{i}$ by minimizing the following contrastive loss, enabling the surrogate model $M_{i}$ to approximate the internal mechanisms of the black-box RAG retrieval model.:
(1) $L=-\frac{1}{|Q|}\sum_{q\in Q}\log\left(\frac{R_{i}(q,d_{+})}{R_{i}(q,d_{+})+\sum R_{i}(q,d_{-})}\right)$
3.4. Opinion Manipulation Attack
After obtaining the surrogate model $M_{i}$, we transform the manipulation of RAG-generated opinions in a black-box scenario into manipulation in a white-box setting. Since we have all the knowledge of the white-box surrogate model $M_{i}$, we directly implement adversarial retrieval attacks on it, generating the adversarial trigger $p_{\text{adv}}$ for the candidate document $d_{t}$ with opinion $S_{t}$. We employ Pairwise Anchor-based Trigger (PAT) (Liu et al., 2023b) strategy for adversarial retrieval attacks, which is commonly used as the baseline in related research. After adding the generated adversarial trigger to the target document with $S_{t}$, the attacker can corrupt the corpus of RAG by inserting the modified candidate document into the RAG corpus.
PAT, as a representative adversarial retrieval attack strategy, adopts a pairwise generation paradigm. Given the target query, the target candidate item, and the top candidate item(named anchor, which is to guide the adversarial text generation), the strategy utilizes gradient optimization of pairwise loss, which is calculated from the candidate item and the anchor, to find the appropriate representation of an adversarial trigger. The strategy also adds fluency constraint and next sentence prediction constraint. By beam search for the words, the final adversarial trigger, denoted as $T_{pat}$, is iteratively generated in an auto-regressive way. We use $T_{\text{pat}}$ as $p_{\text{adv}}$, with its generated optimization function being (Liu et al., 2023b):
(2) $\max\{M_{i}(q,T_{\text{pat}};w)+\lambda_{1}\cdot\log P_{g}(T_{\text{pat}};w)\\ +\lambda_{2}\cdot f_{\text{nsp}}(d_{t},T_{\text{pat}};w)\}$
@@ 3.4. Opinion Manipulation Attack
Subsequently, we insert $p_{\text{adv}}$ into the target document to execute the poisoning attack on the RAG corpus. Our objective of RAG opinion manipulation attack is as follows:
(3) $\max\dfrac{1}{|Q|}\sum_{q\in Q}\text{I}(\text{S}(LLM(q,RM_{k}(q)))=S_{t})$
(4) $\begin{split}s.t.RM(q)&=E(q)\cdot E(d),d\in D(d;d_{t}\oplus p_{\text{adv}})\\ RM_{k}(q)&=\{d_{1},d_{2},...d_{k}\mid argmax_{d\in D}RM(q)\}\end{split}$
@@
4. Experiments
4.1. Dataset
The database of the black-box RAG system in our study consists of MS MARCO Passages Ranking dataset (Nguyen et al., 2016) and PROCON.ORG data [^1] https://testing-www.procon.org/, containing about 500 queries and 500,000 documents. We utilize queries in MSMARCO to obtain the source data for guiding the black-box RAG to generate relevant passages (Zhong et al., 2023) to sample data pairs to train the surrogate model. We use topic queries in PROCON.ORG to get controversial topics arguments for opinion manipulation.
The controversial topic dataset includes over 80 topics, covering fields such as society, health, government, and education. Each controversial topic is discussed from two stances (Pro and Con), with an average of 30 related passages, each holding a certain opinion with stance Pro or Con. We randomly select 30 public controversial topics to conduct opinion manipulation experiment on RAG.
4.2. Experiment Details
@@
We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain [^2] https://python.langchain.com/v0.2/docs/tutorials/qa_chat_history/. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. The LLMs adopted are the open-source models Meta-Llama-3-8B-Instruct (Llama3), Vicuna-13b-v1.5 (Vicuna) and Mixtral-8x7B-Instruct-v0.1 (Mixtral). They perform well across various tasks among all open-source LLMs. The prompt in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain.
@@
We select three representative dense retrieval models, Contriever, Co-Condenser and ANCE, as the target retrieval model (Gao and Callan, 2021; Izacard et al., 2021; Xiong et al., 2020). By convention, we use dot product between the embedding vectors of queries and candidate documents as their similarity score $R$ in the ranking. The surrogate model chosen in this paper is Nboost/pt-bert-base-uncased-msmarco, which is type of BERT and is specifically trained on the MS Marco Passage Ranking dataset.
@@
For a controversial topic $q$, documents $d_{t}$ with the target opinion $S_{t}$ are manipulated by adding adversarial triggers $p_{adv}$ at the beginning. This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3. In this experiment, the number of poisoned documents for FlippedRAG ranges from 5 to 10.
@@
We collect imitation training data in a pairwise way, sampling both positive samples and negative samples. It is worth mentioning that we randomly sample negative examples from the top of the ranking produced by the original surrogate model. We discover that taking highly relevant documents ranked by the surrogate model as negative examples helps the surrogate model learn more features of the target black-box retrieval model by contrastive learning between positive and negative examples.